OpenStudy (anonymous):

The difference of 2 numbers is 38 one number is only 1/3 of the other. what are the two numbers? Answer pleaseee .

OpenStudy (ikram002p):

assume they are x , y now form two equation from the given info. can u do that ??

OpenStudy (anonymous):

@ikram002p i think the two numbers are 57 and 19, but i don't know what is the equation. Can you give me the equation?? please.

OpenStudy (ikram002p):

aha ok x-y=38 y=1/3 x

OpenStudy (anonymous):

@ikram002p my answer is correct?

OpenStudy (ikram002p):

yes its correct, but did u use the equation ir just gussing ??

OpenStudy (anonymous):

trial and error.

OpenStudy (ikram002p):

mmm that is not good... in this case the num is simple , u must try solving it k ??

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

Thanks .

OpenStudy (ikram002p):

wlc

OpenStudy (anonymous):

I have done the equation,it is 57-19=38 one